# A Picture Is Worth a Graph: Blueprint Debate on Graph for Multimodal Reasoning

### v1

@@ Abstract.
This paper presents a pilot study aimed at introducing multi-agent debate into multimodal reasoning. The study addresses two key challenges: the trivialization of opinions resulting from excessive summarization and the diversion of focus caused by distractor concepts introduced from images. These challenges stem from the inductive (bottom-up) nature of existing debating schemes. To address the issue, we propose a deductive (top-down) debating approach called Blueprint Debate on Graphs (BDoG). In BDoG, debates are confined to a blueprint graph to prevent opinion trivialization through world-level summarization. Moreover, by storing evidence in branches within the graph, BDoG mitigates distractions caused by frequent but irrelevant concepts. Extensive experiments validate BDoG, achieving state-of-the-art results in Science QA and MMBench with significant improvements over previous methods.
@@ 1. Introduction
Multimodal reasoning depends on two key aspects: the creation of a unified representation of semantics from different modalities and the integration of these diverse semantics while ensuring logical consistency. While the advancement in large language models (LLMs) has made it possible to represent the semantics in natural languages (Achiam et al., 2023), the integration of diverse semantics remains a challenging issue, even in exclusive NLP tasks. One approach to tackle this challenge is multi-agent debate (MAD), where multiple LLMs act as agents, each contributing their own perspectives on the target topic and reaching a consensus through debates (Liang et al., 2023; Chan et al., 2023). This scheme could be adopted by incorporating a specific LLM for each modality as an agent.
Despite being relatively unexplored in the multimodal domain, MAD encounters numerous challenges in a broader context. MAD may suffer from the trivialization of opinions, which is a result of the summarization step performed at the conclusion of each debating round. The objective of this step is to seek agreement among the participating agents regarding their opinions. Consequently, this process can lead to the debate’s focus being directed towards a general concept, serving as an adaptation to accommodate the diverse range of semantics. One example can be observed in the reasoning of the Multimodal Language Model (MLLM) depicted in Figure 1, where the image modality presents a diverse range of semantics, including bear sedge, earthworm, collared lemming, and others. As a consequence, this can result in the context and summary being trivialized, shifting the emphasis from lichen to a more generalized concept of the tundra ecosystem, wherein both bilberry and mushroom exhibit a high degree of correlation. Similar issue exists when MAD is employed, where the summarizer concludes the diverse semantics into general words like ecosystem and food web, making the conclusion less specific. In addition, MAD may encounter the issue of focus diversion, which occurs when Chain-of-Thoughts (CoT) is utilized and new concepts introduced are highly correlated with a particular concepts (e.g., mathematical model (Cobbe et al., 2021)), leading to an increased weighting of that concept within the context.
@@
To address this issue, we propose an deductive reasoning scheme called Blueprint Debate on Graph (BDoG, pronounced bee-dog). BDoG is inspired by the blueprint debate that has been widely employed in real-world debates, which distinguishes itself from other debates by its concentration on evaluating and refining a proposal (e.g., blueprint) to address specific challenges or issues. BDoG begins by aggregating concepts from modalities and incorporating with their relationships into a initial graph. This graph serves as a blueprint that confines the scope of the discussion rather than having it open to irrelevant semantics as in existing schemes. More importantly, BDoG conducts the debate in a top-down manner by marking down conclusions on the graph. This prevents trivialization as specific concepts are preserved rather than merged into general ones. This can be found from the example shown in Figure 1, where the scope is limited to the tundra ecosystem while specific concepts like mushroom and lichen are retained. Furthermore, the graph provide a compact and high-level guidance for the discussion process. The newly introduced concepts are incorporated into relevant branches instead of remaining as a word-level thoughts within the context. This reduces the likelihood of focus diversion since, in DBoG, the competition of semantics occurs at the branch level rather than the word level. This can be seen from Figure 1, where the most relevant branches related to the soil and caribou standout from the competition, eliminating the irrelevant semantcis effectively. In addition to the advantages of scope-confined guidance and branch-level competition, BDoG also increases explainability, allowing for the tracking of discussion progress. This is evident in Figure 1, in which the progress of debating is more understandable than word-level reasoning (i.e., MLLM-CoT or MAD).
@@ 2.1. Multimodal Reasoning
Multimodal reasoning is a crucial component in the development of advanced artificial intelligence (AI) systems that aim to replicate human-like intelligence (Lu et al., 2022). This type of reasoning enables AI systems to process and analyze information from various sources and forms, such as text, images, audio, and video, in a integrated and coordinated manner (Nam et al., 2017; Cadene et al., 2019). The latest advancements in multimodal large language models, such as BLIP2 (Li et al., 2023a), KOSMOS (Huang et al., 2024) and LLaVA (Liu et al., 2024) have demonstrated significant progress in complex reasoning, as these models (Zhang et al., 2023b) now have the capability to generate step-by-step rationales prior to producing the final answer, following a chain-of-thought (CoT) manner. Zheng et al. (Zheng et al., 2023b) propose a duty-distinct prompting method wherein questions are decomposed into sub-questions to enable deep-layer reasoning. SCITUNE (Horawalavithana et al., 2023) and T-SciQ (Wang et al., 2023b) aim to teach large language models to answer science questions via the generation of mixed rationales derived from both large pretrained models and human annotators. Chameleon (Lu et al., 2024) accomplishes complex multimodal reasoning tasks by integrating various external tools (e.g., large language models, off-the-shelf vision models, and web search engines). Nevertheless, existing methods exhibit limitations as they heavily rely on either few-shot learning or supervision to “guide” the reasoning process. In order to overcome this dependency, we propose the incorporation of debating features into our method. This enables the agents to engage in an adversarial discussion, allowing them to “figure out” the correct direction autonomously. As a result, our approach makes zero-shot learning feasible by reducing the reliance on external guidance or supervision.
@@ 2.3. Graph-augmented LLMs
Prior research has investigated the integration of structured graphs, such as knowledge graphs (KGs), into large language models (LLMs) by embedding the knowledge into the underlying neural networks (Lin et al., 2019; Wang et al., 2023d). Nevertheless, embedding KGs within LLMs may compromise the inherent explainability and adaptability associated with knowledge reasoning and updating (Hu et al., 2023). To tackle these challenges, recent studies have put forth innovative solutions. Li et al. (Li et al., 2023b) proposed an adaptive query generator, facilitating the creation of queries across various query languages (e.g., SPARQL) to infer rationales. Wang et al. (Wang et al., 2023a) devised a structured multi-round question-answering (QA) format, which extracts external knowledge and generates coherent reasoning traces grounded in precise answers. Sun et al. (Sun et al., 2023) introduced Think-on-Graph (ToG), a method that sequentially reasons over KGs to locate relevant triples, thereby supporting the LLM in predicting the final answer. In the context of multimodal reasoning, CCoT (Mitra et al., 2023) substitutes the rationale generation process with scene graph extraction to enhance the compositional capabilities of large multimodal models. KAM-CoT (Mondal et al., 2024), on the other hand, incorporates external KGs during the two-stage training process, yielding state-of-the-art fine-tuning outcomes in multimodal reasoning. In contrast to existing methods that utilize static graphs, our proposed BDoG preserves the dynamics and precision of KGs through iterative updates of entities, attributes, and relationships, guided by a blueprint debate process.
@@ 3. Preliminary
We begin by outlining existing approaches for tackling the multimodal reasoning problem. Figure 2 shows the specific distinct among them. Formally, given a question consisting of tokens, our goal is to identify the correct answer from a set of candidate answers. In the context of multimodal reasoning, the expected answer is intended to be inferred based on a visual context and a textual clue , in addition to the question itself.
@@
DDCoT. The Duty-Distinct Chain of Thought framework proposes a novel approach for deconstructing questions into fundamental sub-questions, similar to breaking down reasoning into elementary steps. Contrary to prior work on conversational agents, Zheng et al. (Zheng et al., 2023b) employ the instruction please decompose the question sequentially into necessary sub-inquiries to acquire the sub-question sequence in a single interaction. Within this framework, the final response is obtained by aggregating the answers to each sub-question and the generated CoT rationale .
@@
MAD. Multi-agent Debate (Liang et al., 2023) presents a promising framework that fosters discursive exchange and cross-pollination of ideas between conversational models. Consider a debate comprising rounds amongst a set of large language models acting as interlocutors. In each round, The proponent generates a rationale and response revised in light of the rationales presented by the opponent in prior turns.
@@ 4. Blueprint Debate on Graph
In this section, we introduce Blueprint Debate-on-Graph (BDoG). As illustrated in Figure 2, BDoG takes a deductive approach instead of inducing answers from word-level thoughts. It utilizes graphs to structure the opinions and proposals provided by agents. This graph-level structuring of the debating context helps minimize opinion trivialization and focus diversion. Furthermore, BDoG adopts an top-down approach which improves multimodal reasoning by iteratively refining an initial proposal, represented as a blueprint graph. This integrates opinions from diverse perspectives through the competition and cooperation among multiple agents.
The debating progress of BDoG at the round can be formulated as a quadruple
where, given a multimodal source set , the debating is conducted among a set of agents , in which each agent can use operations from the set of to propose/summarize opinions by refining the graph-of-thoughts . At the end of the round, is updated to to initiate the next round.
@@ 4.1. Blueprint Initialization
where denotes the union of two sets of graphs. The 2 constraints are as follows: 1) Size Constraint: The size of needs to be restricted within a specific range to prevent an excessive number of clues that could distract the inference or an insufficient number to answer the question effectively. 2) Relevance Constraint: We should merge the relationships extracted from and towards those of the question , ensureing all the knowledge encapsulated in is relevant to the question.
Extensive libraries are available for and , as they have been extensively researched (e.g., named entity recognition, relation extraction for , image captioning, visual grounding for ). However, the recent advancements in multimodal large language models (MLLM) have made it convenient to implement these sub-functions using in-context learning based prompts. For example, to extend the query in the context, we can employ CoT to implement as : Given the question {Q}, please provide the necessary steps to answer this question. where the { } denotes the placeholder in the prompt.
For , its implementation varies depending on LLMs used. For GPT-4, the image needs to be encoded in Base64 format. Gemini utilizes PIL for image encoding. InstructBLIP offers its EVA-G encoder to convert the image into an eigenvector. The can then be implemented as : Given the image {} and question {}, generate a scene graph with evidence to answer the question. Please ensure adherence to following constrains: {}, {}. where two exemplar constraints are The graph must not be empty. Please restrict the maximum number of objects in the graph to 20. The objects and relations within the graph should be pertinent to addressing the question.
@@ 4.2. Agents and Roles
An exemplar implementation is : As {}, you are assigned as an affirmative debater and have been provided with an evidence graph {} for answering the question {} related to the image {}. Try to enhance the graph by incorporating your insights towards an optimal solution. Please ensure adherence to following constrains: {}, {}, {}. Note that we have incorporated the conclusion from (Du et al., 2023; Yin et al., 2023) that the agent’s understanding of the role can be improved by using the {} for targeted personality injection. Furthermore, the personality can be tailored to be specific, such as “Ben, a high school student with an impressive academic record and respected by peers for your knowledge and logical thinking.” The Proponent debate adheres to the Size and Relevance constraints defined in Eq. (3), and it also includes the Compactness Constraint: The refined graph should be as concise as possible, ensuring that the blueprint remains focused.
@@
An exemplar implementation is : As {}, you are assigned as a negative debater and have been provided with an affirmative evidence graph {} for answering the question {} regarding the image {}. Try to detect potential flaws and drawbacks of the graph and update it with your insights. Please ensure adherence to following constrains: {}, {}, {}. The utilization of the functions and fosters an adversarial dynamic between the Proponent and Opponent, ensuring a diverse and comprehensive discussion.
To facilitate the debating, Moderator agents synthesizes the arguments and opinions presented by both the proponent and opponent by merging the and into a conclusion graph as
An exemplar implementation is : As {}, you are assigned as a moderator in a debate and have been provided with an affirmative evidence graph {} and a negative evidence graph {} to address the question {} regarding the image {}. Try to consolidate the two graphs into a single graph towards the optimal solution, and provide a conclusive answer to the question.
@@ 4.3. Debate Progress and Graph Condensation
Initialization and Role Assignment: Once the blueprint has been initialized, the debate commences with the assignment of roles to agents in . Denote the assignment of a role to an agent as , to ensure a balanced debate, an equal number of agents are assigned as Proponents and Opponents, with only one agent assigned as the Moderator. The Role Assignment Regulation is written
@@ 5.1. Backbone Models
To evaluate its performance and generalizability, we have implemented Blueprint Debate-on-Graph (BDoG) using different prevalent multimodal large language models as backbones, including 1) GeminiProVision (Team et al., 2023), an extensively parameterized model developed by Google, 2) InstructBLIP (Dai et al., 2023), which possesses more constrained dimensions and computational resources relative to alternative architectures, and 3) GPT-4 which is the fourth iteration of the GPT model developed by OpenAI. More implementation details can be found from the Appendix.

### v2

# A Picture Is Worth a Graph: A Blueprint Debate Paradigm for Multimodal Reasoning

@@ Abstract.
This paper presents a pilot study aimed at introducing multi-agent debate into multimodal reasoning. The study addresses two key challenges: the trivialization of opinions resulting from excessive summarization and the diversion of focus caused by distractor concepts introduced from images. These challenges stem from the inductive (bottom-up) nature of existing debating schemes. To address the issue, we propose a deductive (top-down) debating approach called Blueprint Debate on Graphs (BDoG). In BDoG, debates are confined to a blueprint graph to prevent opinion trivialization through world-level summarization. Moreover, by storing evidence in branches within the graph, BDoG mitigates distractions caused by frequent but irrelevant concepts. Extensive experiments validate that BDoG is able to achieve state-of-the-art results in ScienceQA and MMBench with significant improvements over previous methods. The source code can be accessed at https://github.com/thecharm/BDoG.
@@ 1. Introduction
Multimodal reasoning depends on two key aspects: the creation of a unified representation of semantics from different modalities and the integration of these diverse semantics while ensuring logical consistency. While the advancement in large language models (LLMs) has made it possible to represent the semantics in natural languages (Achiam et al., 2023; Team et al., 2023), the integration of diverse semantics remains a challenging issue, even in exclusive NLP tasks. One approach to tackle this challenge is multi-agent debate (MAD), where multiple LLMs act as agents, each contributing their own perspectives on the target topic and reaching a consensus through debates (Liang et al., 2023; Chan et al., 2023). This scheme could be adopted by incorporating a specific LLM for each modality as an agent.
While being relatively unexplored in the multimodal domain, MAD encounters numerous challenges in a broader context. It may suffer from the trivialization of opinions, resulting from the summarization step performed at the conclusion of each debating round. The objective of this step is to seek agreement among the participating agents regarding their opinions. Consequently, this process can lead to the debate’s focus being directed towards a general concept, serving as an adaptation to accommodate the diverse range of semantics. One example can be observed in the reasoning of the Multimodal Large Language Model (MLLM) depicted in Figure 1, where the image modality presents a diverse range of semantics, including bear sedge, earthworm, collared lemming, and others. As a consequence, this can result in the context and summary being trivialized, shifting the emphasis from lichen to a more generalized concept of the tundra ecosystem, wherein both bilberry and mushroom exhibit a high degree of correlation. Similar issue exists when MAD is employed, where the summarizer concludes the diverse semantics into general words like ecosystem and food web, making the conclusion less specific. In addition, MAD may encounter the issue of focus diversion, which occurs when Chain-of-Thoughts (CoT) is utilized and new concepts introduced are highly correlated with a particular concepts (e.g., mathematical model (Cobbe et al., 2021)), leading to an increased weighting of that concept within the context.
@@
To address these issues, we propose an deductive reasoning scheme called Blueprint Debate on Graph (BDoG, pronounced bee-dog). BDoG is inspired by the blueprint debate that has been employed in real-world debates, which distinguishes itself from other debates by its concentration on evaluating and refining a proposal (e.g., blueprint) to address specific challenges or issues. BDoG begins by aggregating concepts from modalities and incorporating with their relationships into an initial graph. This graph serves as a blueprint that confines the scope of the discussion rather than having it open to irrelevant semantics as in existing schemes. More importantly, BDoG conducts the debate in a top-down manner by marking down conclusions on the graph. This prevents trivialization as specific concepts are preserved rather than merged into general ones. This can be found from the example shown in Figure 1, where the scope is limited to the tundra ecosystem while specific concepts like mushroom and lichen are retained. Furthermore, the graph provide a compact and high-level guidance for the discussion process. The newly introduced concepts are incorporated into relevant branches instead of remaining as a word-level thoughts within the context. This reduces the likelihood of focus diversion since, in BDoG, the competition of semantics occurs at the branch level rather than the word level. This can be seen from Figure 1, where the most relevant branches related to the soil and caribou standout from the competition, eliminating the irrelevant semantics effectively. In addition to the advantages of scope-confined guidance and branch-level competition, BDoG also increases explainability, allowing for the tracking of discussion progress (Figure 1).
@@ 2.1. Multimodal Reasoning
Multimodal reasoning is a crucial component in the development of advanced artificial intelligence (AI) systems that aim to replicate human-like intelligence (Lu et al., 2022; lv2023duet). The latest advancements in multimodal large language models, such as BLIP2 (Li et al., 2023b) and LLaVA (Liu et al., 2024) have demonstrated significant progress in complex reasoning, as these models (Zhang et al., 2023b) now have the capability to generate step-by-step rationales prior to producing the final answer, following a chain-of-thought (CoT) manner. Zheng et al. (Zheng et al., 2023b) propose a duty-distinct prompting method wherein questions are decomposed into sub-questions to enable deep-layer reasoning. SCITUNE (Horawalavithana et al., 2023) and T-SciQ (Wang et al., 2023b) aim to teach large language models to answer science questions via the generation of mixed rationales derived from both large pretrained models and human annotators. Chameleon (Lu et al., 2024) accomplishes complex multimodal reasoning tasks by integrating various external tools (e.g., large language models, off-the-shelf vision models, and web search engines).
@@ 2.3. Graph-augmented LLMs
Prior research has investigated the integration of structured graphs, such as knowledge graphs (KGs), into large language models (LLMs) by embedding the knowledge into the underlying neural networks (wei2008fusing; Wang et al., 2023d). Nevertheless, embedding KGs within LLMs may compromise the inherent explainability and adaptability associated with knowledge reasoning and updating (Hu et al., 2023). To tackle these challenges, recent studies have put forth innovative solutions. Li et al. (Li et al., 2023c) propose an adaptive query generator, facilitating the creation of queries across various query languages (e.g., SPARQL) to infer rationales. Wang et al. (Wang et al., 2023a) devise a structured multi-round question-answering (QA) format, which extracts external knowledge and generates coherent reasoning traces grounded in precise answers. Sun et al. (Sun et al., 2023) introduce Think-on-Graph (ToG), a method that sequentially reasons over KGs to locate relevant triples, thereby supporting the LLM in predicting the final answer. In the context of multimodal reasoning, CCoT (Mitra et al., 2023) substitutes the rationale generation process with scene graph extraction to enhance the compositional capabilities of large multimodal models. KAM-CoT (Mondal et al., 2024), on the other hand, incorporates external KGs during the two-stage training process, yielding state-of-the-art fine-tuning outcomes in multimodal reasoning. In contrast to existing methods that utilize static graphs, our proposed BDoG preserves the dynamics and precision of KGs through iterative updates of entities, attributes, and relationships, guided by a blueprint debate process.
@@ 3. Preliminary
We begin by outlining existing approaches for tackling the multimodal reasoning problem. Figure 2 shows the specific distinction among them. Formally, given a question consisting of tokens, our goal is to identify the correct answer from a set of candidate answers. In the context of multimodal reasoning, the expected answer is intended to be inferred based on a visual context and a textual clue , in addition to the question itself.
@@
DDCoT. The Duty-Distinct Chain of Thought framework proposes a novel approach for deconstructing questions into fundamental sub-questions, similar to breaking down reasoning into elementary steps. Contrary to prior work on conversational agents, Zheng et al. (Zheng et al., 2023b) employ the instruction to acquire the sub-question sequence in a single interaction. Within this framework, the final response is obtained by aggregating the answers to each sub-question and the generated CoT rationale .
@@
MAD. MAD (Liang et al., 2023) presents a promising framework that fosters discursive exchange and cross-pollination of ideas between conversational models. Consider a debate comprising rounds amongst a set of large language models acting as interlocutors, the proponent generates a rationale and response revised in the light of rationales presented by the opponent in prior turns.
@@ 4. Blueprint Debate on Graph
In this section, we introduce Blueprint Debate-on-Graph (BDoG). As illustrated in Figure 2, BDoG takes a deductive approach instead of inducing answers from word-level thoughts. It utilizes graphs to structure the opinions and proposals provided by agents. This graph-level structuring of the debating context helps to minimize opinion trivialization and focus diversion. Furthermore, BDoG adopts a top-down approach which improves multimodal reasoning by iteratively refining an initial proposal, represented as a blueprint graph. This integrates opinions from diverse perspectives through the competition and cooperation among multiple agents.
The BDoG at the round can be formulated as a quadruple
where, given a multimodal source set , the debating is conducted among a set of agents , in which each agent uses operations from the set to propose opinions by refining the graph-of-thoughts . At the end of the round, is updated to to initiate the next round.
@@ 4.1. Blueprint Initialization
where denotes the union of two sets of graphs. The 2 constraints are as follows: 1) Size Constraint: The size of needs to be restricted within a specific range to prevent an excessive number of clues that could distract the inference or an insufficient number to answer the question effectively. 2) Relevance Constraint: We should merge the relationships extracted from and towards those of the question , ensuring all the knowledge encapsulated in is relevant to the question. Extensive libraries are available for and , as they have been extensively researched (e.g., named entity recognition (Wu et al., 2020), relation extraction (Zheng et al., 2021) for , image captioning (Zhao and Wu, 2023), visual grounding (Li et al., 2023a) for ). However, the recent advancements in multimodal large language models (MLLM) have made it convenient to implement these sub-functions using in-context learning based prompts. For example, to extend the query in the context, we can employ CoT to implement as {mdframed}[hidealllines=true,backgroundcolor=gray!20]
: Given the question {Q}, please provide the necessary steps to answer this question. where the { } denotes the placeholder in the prompt.
For , its implementation varies depending on LLMs used. For GPT-4, the image needs to be encoded in Base64 format. Gemini utilizes PIL for image encoding. InstructBLIP offers its EVA-G encoder to convert the image into an eigenvector. The can then be implemented as {mdframed}[hidealllines=true,backgroundcolor=gray!20]
: Given the image {} and question {}, generate a scene graph with evidence to answer the question. Please ensure adherence to following constrains: {}, {}. where two exemplar constraints are {mdframed}[hidealllines=true,backgroundcolor=gray!20]
The graph must not be empty. Please restrict the maximum number of objects in the graph to 20. {mdframed}[hidealllines=true,backgroundcolor=gray!20]
The objects and relations within the graph should be pertinent to addressing the question.
@@ 4.2. Agents and Roles
An exemplar implementation is {mdframed}[hidealllines=true,backgroundcolor=gray!20]
: As {}, you are assigned as an affirmative debater and have been provided with an evidence graph {} for answering the question {} related to the image {}. Try to enhance the graph by incorporating your insights towards an optimal solution. Please ensure adherence to following constrains: {}, {}, {}. Note that we have incorporated the conclusion from (Du et al., 2023; Yin et al., 2023) that the agent’s understanding of the role can be improved by using the {} for targeted personality injection. Furthermore, the personality can be tailored to be specific, such as “Ben, a high school student with an impressive academic record and respected by peers for your knowledge and logical thinking.” The Proponent debate adheres to the Size and Relevance constraints defined in Eq. (3), and it also includes the Compactness Constraint: The refined graph should be as concise as possible, ensuring that the blueprint remains focused.
@@
An exemplar implementation is {mdframed}[hidealllines=true,backgroundcolor=gray!20]
: As {}, you are assigned as a negative debater and have been provided with an affirmative evidence graph {} for answering the question {} regarding the image {}. Try to detect potential flaws and drawbacks of the graph and update it with your insights. Please ensure adherence to following constrains: {}, {}, {}. The utilization of the functions and fosters an adversarial dynamic between the Proponent and Opponent, ensuring a diverse and comprehensive discussion.
To facilitate the debating, Moderator agents synthesize the arguments and opinions presented by both the proponent and opponent by merging the and into a conclusion as
An exemplar implementation is {mdframed}[hidealllines=true,backgroundcolor=gray!20]
: As {}, you are assigned as a moderator in a debate and have been provided with an affirmative evidence graph {} and a negative evidence graph {} to address the question {} regarding the image {}. Try to consolidate the two graphs into a single graph towards the optimal solution, and provide a conclusive answer to the question.
@@ 4.3. Debate Progress and Graph Condensation
Initialization and Role Assignment: Once the blueprint has been initialized, the debate commences with the assignment of roles to agents in . Denote the assignment of a role to an agent as , to ensure a balanced debate, an equal number of agents are assigned as Proponents and Opponents, with only one agent assigned as the Moderator. The Role Assignment Regulation is
@@ 5.1. Backbone Models
To evaluate its performance and generalizability, we have implemented Blueprint Debate-on-Graph (BDoG) using different prevalent multimodal large language models as backbones, including 1) GeminiProVision (Team et al., 2023), an extensively parameterized model developed by Google, 2) InstructBLIP (Dai et al., 2023) and LLaVA-v1.5 (Liu et al., 2024), which possesses more constrained dimensions and computational resources relative to alternative architectures, and 3) GPT-4 (Achiam et al., 2023) which is the fourth iteration of the GPT model developed by OpenAI. More implementation details can be found from the Appendix.